OpenStudy (anonymous):

Parker earned $32 by selling 27 glasses of fresh-squeezed lemonade at his lemonade stand on a Saturday morning. He charged $1.00 for a small glass and $1.50 for a large glass. Use a model to determine how many large glasses he sold. I just need the answer.

OpenStudy (anonymous):

$1 * x = number of glasses sold $0.50 * y = additional charge for large glass. So $1 * x +$0.50 *y = $32 x = 27 Can you take it from here?